OpenStudy (princess_kylee):

The length of a side in an equilateral triangle is 5 inches. Find the area of the triangle.

OpenStudy (wolf1728):

Triangle Area = (1/2) base * height

OpenStudy (wolf1728):

The base is 5 inches the height has to be calculated

OpenStudy (wcrmelissa2001):

find the height using pythagoras' theorem.  So to find the height, it's sqrt(5^2-2.5^2)=4.33 (3 s.f.)